OpenStudy (anonymous):

Find this number to the nearest tenth. What number is 11 1/3 % of 215?

OpenStudy (anonymous):

convert your percentage to a decimal number. but first, interpret 1/3 as a decimal number first. 1/3 is approximately 0.3333 (continuous decimal) so 11 1/3% is actually 11.3333% now convert 11.33% into decimal by dividing 11.3333 by 100 =11.3333/100 =0.113333 Then now we can multiply it by 215 to get the answer: = 0.113333*215 =24.366666667 The answer has a continuous decimal value so cut and round the answer depending on how much decimal number will you need to have to get the right answer.

OpenStudy (anonymous):

I think I get it. Thanks. :)